Find the EXACT VALUE of the following, A) (SIN 30°)²+cot60+(csc45)² B) (csc60)/sin45° *for letter B, it is (csc60) over sin45°
Found 3 solutions by MowMow, MathTherapy, ikleyn:
Answer by MowMow(42)

A) (SIN 30°)²+cot60+(csc45)² = (.5)^2 + .5774 + 1.414^2 = .25+.577+2 = 2.8274
cot60 = 1/Tan60
csc45 = 1/Sin45
B) (csc60)/sin45° = (1/Sin60)/Sin45 = 1.1547/.7071 = 1.633

Answer by MathTherapy(10555)


Find the EXACT VALUE of the following, A) (SIN 30°)²+cot60+(csc45)² B) (csc60)/sin45° *for letter B, it is (csc60) over sin45°
If you don't know by now, that person DIDN'T give you EXACT values. 


Answer by ikleyn(52835)

.

@MowMow, it is clear from your post, that you don't know basics of Trigonometry.


I observed your contribution to the forum,

and with 179% confidence I can say that you have no necessary knowledge to be a Math tutor.

I ask you STOP posting to this forum.